OpenStudy (anonymous):

please help me Find the Geometric Mean of the pair of numbers 3 and 12. * Remember the Geometric Mean of two numbers a and b is found by taking a/x = x/b, where x is the Geometric Mean. Solve for x

OpenStudy (anonymous):



OpenStudy (johnweldon1993):

Oh okay if we have to do it that way...we have \[\large \frac{3}{x} =\frac{x}{12}\] Cross multiply \[\large x^2 = 36\] \[\large x = \sqrt{36}\] \[\large x = 6\]